### v1

# Controllable Graph Generation with Diffusion Models via Inference-Time Tree Search Guidance

Abstract.
Graph generation is a fundamental problem in graph learning with broad applications across Web-scale systems, knowledge graphs, and scientific domains such as drug and material discovery. Recent approaches leverage diffusion models for step-by-step generation, yet unconditional diffusion offers little control over desired properties, often leading to unstable quality and difficulty in incorporating new objectives. Inference-time guidance methods mitigate these issues by adjusting the sampling process without retraining, but they remain inherently local, heuristic, and limited in controllability. To overcome these limitations, we propose TreeDiff, a Monte Carlo Tree Search (MCTS) guided dual-space diffusion framework for controllable graph generation. TreeDiff is a plug-and-play inference-time method that expands the search space while keeping computation tractable. Specifically, TreeDiff introduces three key designs to make it practical and scalable: (1) a macro-step expansion strategy that groups multiple denoising updates into a single transition, reducing tree depth and enabling long-horizon exploration; (2) a dual-space denoising mechanism that couples efficient latent-space denoising with lightweight discrete correction in graph space, ensuring both scalability and structural fidelity; and (3) a dual-space verifier that predicts long-term rewards from partially denoised graphs, enabling early value estimation and removing the need for full rollouts. Extensive experiments on 2D and 3D molecular generation benchmarks, under both unconditional and conditional settings, demonstrate that TreeDiff achieves state-of-the-art performance. Notably, TreeDiff exhibits favorable inference-time scaling: it continues to improve with additional computation, while existing inference-time methods plateau early under limited resources.
Graph Generation, Diffusion Models, Inference-Time Guidance, Monte Carlo Tree Search
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Graphs are powerful structures for modeling complex relationships among entities (Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Wang et al., 2024a, 2025c) and form the backbone of numerous Web-related systems, such as knowledge graphs, recommender networks, and online interaction graphs. The ability to generate and manipulate such graphs under domain-specific constraints has become a fundamental problem in graph learning (Bonifati et al., 2020; Simonovsky and Komodakis, 2018; Shi et al., 2020; Vignac et al., 2023a; Liu et al., 2024). For example, in molecular and material design (Liu et al., 2024)—where molecules can be viewed as structured graphs—generation models aim to discover new candidates that satisfy specific chemical or physical properties. Similarly, in broader Web contexts, graph generation underpins applications such as link prediction (Hamilton et al., 2017), knowledge graph completion (Chen et al., 2020), and the simulation of evolving information networks (Myers et al., 2014). Despite these diverse applications, generating graphs that are valid, diverse, and aligned with complex objectives remains a significant algorithmic challenge.
Figure 1. Inference-time scaling behavior on four generation benchmarks. TreeDiff exhibits consistent gains with additional inference computation, whereas existing approaches—including standard diffusion (Jo et al., 2022), Best-of-N (Ma et al., 2025), and state-of-the-art inference-time guidance (Li et al., 2024)—saturate quickly, highlighting the scalability of our approach.
Various generative models have been explored for graph generation (Jensen, 2019; Brown et al., 2019; Jin et al., 2018; Shi et al., 2020; Zang and Wang, 2020), among which diffusion models (Vignac et al., 2023a; Jo et al., 2022; Hoogeboom et al., 2022; Xu et al., 2023; Liu et al., 2024) have recently gained prominence due to their ability to generate diverse, high-quality samples. Specifically, a diffusion model (Ho et al., 2020) learns the reverse denoising process during training to approximate the data distribution. During inference, it starts from noise and iteratively denoises to recover valid samples, ensuring consistency between training and inference. Despite the success of diffusion models in graph generation, they face two major limitations that hinder their practicality. First, controllability is limited (Peebles and Xie, 2023; Podell et al., 2024): when trained without explicit conditions or constraints, diffusion models primarily approximate the unconditional data distribution, lacking a principled way to enforce target properties during sampling. Second, generation can be unstable (Watson et al., 2023): the stochasticity of noise sampling often leads to invalid or low-quality outputs, reflecting the inherent fragility of the denoising process. These issues are especially severe in graph generation, where tasks like drug discovery demand molecules with domain-specific attributes (e.g., high QED, low SAS), yet such properties are rarely optimized during training. Moreover, graphs are far less tolerant to perturbations than images—even minor structural changes can drastically alter graph semantics (Wang et al., 2024b, 2025b).
To this end, recent works have proposed inference-time guidance methods (Ma et al., 2025; Dorna et al., 2024; Li et al., 2024), which steer the sampling process at test time without altering training. Specifically, these methods refine sampling under given conditions and prune or adjust invalid steps when necessary, thereby enabling more stable and controllable generation. For instance, Ma et al. (2025) maintain a pool of candidate noises and perform zero-order search to expand promising branches during inference for better generation quality. Similarly, Dorna et al. (2024) leverage an auxiliary classifier as a reward model to estimate the quality of intermediate states and inject gradient signals for controllable graph generation. Despite their effectiveness, existing approaches remain fundamentally local: they make decisions based on step-level feedback, greedily correcting or discarding samples according to the current state without anticipating long-term consequences. This reliance on local information limits their ability to plan across long denoising trajectories or to satisfy complex, multi-objective constraints, leaving controllability and stability still restricted (Yang et al., 2023). Empirically, this limitation manifests in poor inference-time scaling: as illustrated in Figure 1, increasing computation does not improve—and can even harm—the performance of existing approaches, underscoring their inability to benefit from additional computing resources.
Figure 2. Comparison among diffusion-based graph generation approaches. (a) Standard graph diffusion iteratively denoises each step without correction. (b) Inference-time guidance introduces local corrections via classifier gradients but remains short-sighted. (c) Monte Carlo Tree Search (MCTS) enables structured lookahead over denoising trajectories, allowing for globally balanced exploration and exploitation.
To overcome the limitation, we propose using Monte Carlo Tree Search (MCTS) (Silver et al., 2016) as a novel inference-time controller for graph generation. Unlike greedy stepwise heuristics, MCTS enables lookahead search, evaluates future trajectories, and integrates multiple objectives in a principled manner, offering a more stable and controllable generation process. However, directly applying MCTS to diffusion-based graph generation is non-trivial and introduces several unique challenges. (1) Tree Depth Explosion: Diffusion models typically require hundreds or even thousands of denoising steps, making one-step-per-node expansion computationally infeasible. This calls for mechanisms that can reduce the effective depth of the search while preserving the overall denoising trajectory. (2) Ill-Defined Node Expansion Space: Defining the expansion space is also problematic. Denoising directly in graph space ensures stability and validity but is extremely slow (Vignac et al., 2023a), while relying solely on embedding space is efficient but unstable (Dorna et al., 2024), as small perturbations can lead to drastic structural changes. This highlights the need for strategies that balance efficiency with structural reliability during node expansion. (3) Costly and Complex Evaluation: MCTS requires reliable evaluation of future outcomes, yet in graph diffusion, full rollouts are prohibitively expensive: one must complete the entire denoising trajectory and compute molecular properties (e.g., QED, SAS), while also enforcing validity constraints. This makes it essential to design evaluation schemes that approximate long-term quality efficiently while pruning infeasible branches early.
To overcome these challenges, we introduce TreeDiff, a MCTS-guided dual-space diffusion framework for controllable graph generation. Specifically, TreeDiff incorporates three key designs: (1) a macro-step expansion strategy that reduces tree depth by grouping multiple denoising steps into one node, (2) a dual-space denoising mechanism that combines efficiency in the embedding space with stability in the graph space, and (3) a dual-space verifier that provides long-term value estimates to guide global search under complex validity constraints. Extensive experiments on 2D and 3D molecular benchmarks show that TreeDiff achieves state-of-the-art performance on both unconditional and conditional generation tasks. Notably, as shown in Figure 1, TreeDiff exhibits favorable inference-time scaling behavior—continuing to improve as more computation is allocated, whereas existing methods quickly plateau.
2. Related Work
Graph Generation. Graph generation (Bonifati et al., 2020; Thompson et al., 2022; Tang et al., 2022) learns distributions over graph topology and attributes, enabling both unconditional and conditional synthesis. However, validity, property control, and efficiency remain difficult to achieve simultaneously (Maziarz et al., 2021; Xu et al., 2021). Prior work mainly falls into four families: autoregressive models allow rule injection but are order-sensitive and slow (You et al., 2018; Liao et al., 2019; Shi et al., 2020); VAE-based methods support latent interpolation but impose size limits (Simonovsky and Komodakis, 2018; Jin et al., 2018); GAN-based approaches provide one-shot fidelity but suffer from instability (De Cao and Kipf, 2018); and flow-based models yield tractable likelihoods but face dequantization bias (Zang and Wang, 2020; Luo et al., 2021). More recently, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020) have emerged as the dominant paradigm, offering high-quality graph generation.
Graph Diffusion Models. Diffusion models (Ho et al., 2020; Rombach et al., 2022; Gong et al., 2022; Nie et al., 2025) have recently been adapted to graphs by injecting noise into discrete topology and attributes and learning a reverse denoising process. Representative works include GDSS (Jo et al., 2022), which learns joint node & adjacency scores in a continuous-time SDE; DiGress (Vignac et al., 2022), which applies discrete transformations over nodes and edges; GeoDiff (Xu et al., 2022), which performs SE(3)-equivariant diffusion for 3D conformations; and NVDiff (Chen et al., 2022), which denoises node embeddings for scalability. Extensions further explore conditional generation (Liu et al., 2023; Jensen, 2019; Xie et al., 2021; Lee et al., 2023; Guan et al., 2023, 2024) and efficient sampling (Luo et al., 2023; Cho et al., 2023). Despite these advances, controllability and stability remain open challenges: enforcing graph-level constraints consistently across multi-step denoising is still unreliable. Our proposed TreeDiff directly addresses these issues at inference time without retraining.
Inference-Time Guidance. For graph diffusion, inference-time guidance (Uehara et al., 2025) offers a way to enforce validity and target properties without retraining, by steering the denoising trajectory at test time. Existing approaches can be grouped into three categories: (i) classifier guidance (Dhariwal and Nichol, 2021; Ho and Salimans, 2022; Kim et al., 2022; Chung et al., 2022; Dorna et al., 2024), which trains auxiliary models to score graph properties (e.g., QED/SA) and provide gradients or rewards; (ii) sampling-based methods (Wu et al., 2023; Cardoso et al., 2023; Phillips et al., 2024; Li et al., 2024; Ma et al., 2025; Nichol et al., 2021), which resample from candidate noises or trajectories and retain the best ones; and (iii) search-based methods (Yoon et al., 2025; Jain et al., 2025; Oshima et al., 2025; Li et al., 2024; Hubert et al., 2021; Feng et al., 2023; Hao et al., 2023; Kajita et al., 2020; Swanson et al., 2024), which keep a small frontier and apply heuristics or light backtracking. While these methods improve controllability, in graph domains they remain limited: feedback is often noisy, validity is fragile, and scaling candidates rarely yields consistent gains. Our proposed TreeDiff instead employs MCTS for global planning, integrating property guidance with validity checks along full trajectories to achieve robust, controllable, and scalable inference.
3. Preliminary
Problem Definition. We study the problem of graph generation. Formally, let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\mathbf{X}},{\mathbf{E}})$ denote a graph, where $\mathcal{V}$} is the set of nodes with $|{\mathcal{V}}|=N$, ${\mathcal{E}}\subseteq{\mathcal{V}}\times{\mathcal{V}}$ is the set of edges with $|{\mathcal{E}}|=E$, and $\mathbf{X}$} and $\mathbf{E}$} represent the node and edge feature matrices, respectively. Each node $v\in{\mathcal{V}}$ is associated with a feature vector $\mathbf{x}_{v}\in\mathbb{R}^{d_{n}}$, and each edge $e\in{\mathcal{E}}$ is associated with a feature vector $\mathbf{e}_{e}\in\mathbb{R}^{d_{e}}$ (if edge attributes are available). Given a target distribution $\mathcal{P}_{\text{data}}$ over graphs, the goal is to learn a generative model $p_{\theta}({\mathcal{G}})$ such that $p_{\theta}({\mathcal{G}})\approx\mathcal{P}_{\text{data}}$. That is, the model should generate graphs that are both structurally valid and semantically consistent with the observed data, while preserving diversity and faithfully capturing the underlying distribution.
Graph Diffusion Model. We adopt the diffusion framework in the context of graph generation. Given a graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\mathbf{X}},{\mathbf{E}})$, the forward process constructs a sequence of latent variables $\{{\mathcal{G}}_{t}\}_{t=0}^{T}$, where ${\mathcal{G}}_{0}={\mathcal{G}}$ and ${\mathcal{G}}_{T}$ approximates a noise distribution. At each step, noise is added according to a predefined Markov transition:
(1) $q({\mathcal{G}}_{t}\mid{\mathcal{G}}_{t-1})=\mathcal{N}\!\left(\sqrt{1-\beta_{t}}\,{\mathcal{G}}_{t-1},\,\beta_{t}{\mathbf{I}}\right),$
where $\{\beta_{t}\}_{t=1}^{T}$ is a variance schedule. The reverse process aims to learn a parameterized distribution $p_{\theta}({\mathcal{G}}_{t-1}\mid{\mathcal{G}}_{t})$ that approximates the true posterior $q({\mathcal{G}}_{t-1}\mid{\mathcal{G}}_{t},{\mathcal{G}}_{0})$. Sampling proceeds by iteratively denoising from Gaussian noise toward the target graph $\mathcal{G}_{0}$. The training objective is typically a reweighted denoising score matching loss:
(2) $\mathcal{L}_{\text{diff}}(\theta)=\mathbb{E}_{t,{\mathcal{G}}_{0},\epsilon}\!\left[\|\epsilon-\epsilon_{\theta}({\mathcal{G}}_{t},t)\|^{2}\right],$
where $\epsilon$ is Gaussian noise and $\epsilon_{\theta}$ is a neural network that predicts the noise component.
In graph generation, diffusion can be applied either directly in the graph space (adjacency and feature matrices) or in a continuous latent space learned by an encoder. Graph-space diffusion preserves structural semantics explicitly, but the discrete and combinatorial nature of graphs makes the denoising process unstable and computationally expensive. Latent-space diffusion alleviates these issues by operating in a continuous space, enabling smoother optimization and more scalable training; however, it often loses fine-grained structural information, leading to limited controllability. This trade-off highlights the need for designs that can better balance structural fidelity and generative flexibility.
Monte Carlo Tree Search. The denoising process in diffusion models can be naturally viewed as a sequential decision problem, where each step refines the current state and local errors may accumulate into invalid or low-quality graphs. This motivates the use of search-based methods that can plan across long denoising trajectories. Among them, MCTS is a widely used algorithm for decision-making in large spaces under uncertainty. Formally, MCTS builds a search tree $\mathcal{T}$}, where each node $s\in{\mathcal{T}}$ represents a state (e.g., a partially denoised graph), and edges correspond to actions $a\in\mathcal{A}(s)$ (e.g., candidate denoising steps). Each iteration of MCTS consists of four phases: (i) Selection, choosing a path from the root to a leaf according to an exploration–exploitation policy (e.g., choosing a partial denoising trajectory); (ii) Expansion, adding a new child node to represent an unexplored state (e.g., the next denoising step); (iii) Simulation, estimating the long-term outcome by rolling out a trajectory from the new node (e.g., sampling a complete graph), yielding a reward $R$; and (iv) Backpropagation, propagating $R$ upward to update the statistics of nodes along the path. By adaptively allocating computation to promising branches, MCTS provides lookahead planning, maintains multiple candidate trajectories, and integrates domain-specific objectives.
4. Methodology
Figure 3. TreeDiff integrates MCTS with diffusion-based graph generation. (a) TreeDiff follows the standard MCTS pipeline but introduces new designs in the Expansion and Simulation phases. (b) Macro-step expansion reduces tree depth by grouping multiple denoising updates. (c) Dual-space denoising couples latent diffusion with discrete structural correction to ensure validity. (d) Dual-space verifier predicts long-term rewards from latent and graph embeddings, enabling rollout-free evaluation.
4.1. Framework
We propose TreeDiff, an MCTS-guided diffusion framework for controllable and stable graph generation. The core idea is to view the denoising process as a sequential decision problem and augment it with structured search, so that alternative denoising trajectories can be explored, compared, and optimized during inference. While MCTS provides a natural mechanism for long-term planning, directly applying it to diffusion sampling is infeasible due to three key obstacles: (i) the enormous depth of diffusion trajectories, (ii) the tension between efficiency in latent space and stability in graph space, and (iii) the prohibitive cost of evaluating complete rollouts.
To address these issues, TreeDiff incorporates three complementary components (Figure 3). First, a macro-step expansion strategy reduces tree depth by grouping multiple denoising steps into a single node transition. Second, a dual-space denoising mechanism couples latent-space denoising with graph-space guidance, balancing efficiency with structural fidelity. Third, a dual-space verifier combines fast validity checks with more expensive property-based objectives, enabling efficient yet reliable reward estimation. These designs allow TreeDiff to allocate inference-time computation adaptively, achieving improvements in graph quality and controllability. The implementation details are provided in Appendix C.
4.2. Macro-Step Expansion
In MCTS, the Expansion phase controls how new nodes are added to the search tree. For diffusion-based graph generation, each reverse step only makes a minor refinement to the current graph state, yielding an extremely deep, fine-grained denoising trajectory. If expansion were performed at every diffusion step, the tree would contain hundreds or thousands of layers with near-duplicate parent/child states, making exploration prohibitively slow and preventing meaningful long-horizon planning. It is therefore crucial to redefine the granularity of expansion so that the tree grows over coarser, semantically meaningful transitions rather than raw step-level perturbations. We introduce a macro-step expansion strategy that groups multiple reverse denoising updates into a single transition. Instead of adding a child for each step $t\!\to\!t-1$, TreeDiff expands from ${\mathcal{G}}_{t}$ directly to ${\mathcal{G}}_{t-k}$ for a chosen horizon $k$, thereby preserving the overall denoising dynamics while drastically reducing the effective search depth. From the MCTS perspective, this turns expansion from fine-grained local moves into a hierarchical exploration of the denoising landscape.
Formally, let ${\mathcal{G}}_{t}$ denote the intermediate graph at reverse step $t$. A single reverse update, induced by the learned noise predictor $\epsilon_{\theta}({\mathcal{G}}_{t},t)$ and variance schedule $\{\beta_{t}\}_{t=1}^{T}$, can be written as a sampler ${\mathcal{G}}_{t-1}\;=\;\Phi_{\theta}({\mathcal{G}}_{t},t;\,\xi_{t})$, where $\Phi_{\theta}$ denotes the one-step reverse transition and $\xi_{t}\sim\mathcal{N}(0,{\mathbf{I}})$ is the step-wise randomness. We define a $k$-step macro-step operator by composing $k$ such updates:
${\mathcal{G}}_{t-k}$ $=\Phi_{\theta}^{(k)}\big({\mathcal{G}}_{t},t;\xi_{t-k+1:t}\big)$
(3) $\triangleq\Phi_{\theta}\big(\Phi_{\theta}^{(k-1)}({\mathcal{G}}_{t},t;\xi_{t-k+2:t}),t-k+1;\xi_{t-k+1}\big),$
with $\xi_{t-k+1:t}=\{\xi_{t-k+1},\ldots,\xi_{t}\}$. During Expansion, TreeDiff samples a small set of candidate macro-steps and inserts the resulting as children of ${\mathcal{G}}_{t}$ in the search tree $\mathcal{T}$}.
Adaptive Step Sampling. To further enhance exploration diversity, TreeDiff dynamically determines the macro-step length $k$ at each expansion rather than fixing it across the tree. For a node corresponding to step $t$, we estimate the remaining denoising steps $T_{\text{rem}}=t$ and the remaining allowable tree depth $D_{\text{rem}}$. A base step size is computed as $k_{\text{base}}=\frac{T_{\text{rem}}}{D_{\text{rem}}}$, representing the average number of diffusion updates each future expansion should roughly cover. Then, for every expansion, TreeDiff samples the actual $k$ from a Gaussian distribution centered at this estimate:
(4) $k\sim\mathcal{N}\!\big(k_{\text{base}},\,(\sigma_{k}\,k_{\text{base}})^{2}\big),$
and clips it to ensure $1\leq k\leq T_{\text{rem}}$, where $\sigma_{k}$ is a deviation coefficient (set to $0.1$ by default). This stochastic resampling introduces controlled randomness in temporal granularity: some branches advance aggressively with longer macro-steps, while others refine locally with shorter ones. When the remaining steps fall below $k_{\text{base}}$, TreeDiff performs the remaining denoising to reach ${\mathcal{G}}_{0}$. This adaptive sampling further balances exploration and exploitation in MCTS, leading to more robust and diverse denoising trajectories.
4.3. Dual-Space Denoising
While macro-step expansion enables TreeDiff to explore diffusion trajectories efficiently, it leaves open a key question for MCTS: in which representation should Selection and Expansion operate? Each node in the search tree corresponds to a partially denoised graph, which can be expressed either in the continuous latent space or in the discrete graph space. Choosing either space alone introduces a fundamental trade-off. Operating purely in the latent space provides smooth, differentiable transitions suitable for selection policies such as UCB (Auer et al., 2002) or PUCT (Silver et al., 2016), but decoded trajectories often drift into invalid graphs, undermining reliable value estimation. Conversely, expanding and evaluating directly in the discrete graph space preserves structural validity and enables domain-specific scoring (e.g., chemical properties), but leads to combinatorial branching and discontinuous transitions, making selection noisy and unstable. This mismatch between continuous exploration and discrete evaluation disrupts MCTS’s ability to balance exploitation and exploration across the denoising trajectory. To bridge this gap, TreeDiff introduces a dual-space denoising mechanism that couples continuous latent diffusion with lightweight discrete correction in graph space.
Specifically, dual-space denoising governs the internal dynamics of each MCTS expansion with $k$ steps by weaving continuous diffusion and discrete refinement into a coherent update rule. Let ${\mathcal{G}}_{t}^{\text{latent}}$ denote the latent representation of a partially denoised graph at reverse step $t$, and ${\mathcal{G}}_{t}^{\text{structure}}$ its decoded discrete counterpart. TreeDiff first advances $n$ latent denoising steps to reach ${\mathcal{G}}_{t-n}^{\text{latent}}$, then projects this intermediate state into the structural domain, ${\mathcal{G}}_{t-n}^{\text{structure}}=\mathrm{Dec}({\mathcal{G}}_{t-n}^{\text{latent}},t-n)$, where a lightweight graph denoiser $p_{\psi}$ performs $m$ categorical refinement steps ($m\ll n$) to enforce topological consistency and inject symbolic priors. The refined graph ${\mathcal{G}}_{t-n-m}^{\text{structure}}$ is re-encoded into a structure-aware latent ${\mathcal{G}}_{t-n-m}^{\text{latent}}=\mathrm{Enc}({\mathcal{G}}_{t-n-m}^{\text{structure}},t-n-m)$, embedding discrete priors back into the continuous manifold.
Rather than treating this reprojected latent as a new trajectory, TreeDiff interprets it as a directional guide that reshapes the remaining diffusion. A guidance vector is computed as
(5) $g_{t-n}=\nabla_{{\mathcal{G}}_{t-n}^{\text{latent}}}\!\left(-\frac{1}{2\sigma_{g}^{2}}\big\|{\mathcal{G}}_{t-n-m}^{\text{latent}}-{\mathcal{G}}_{t-n}^{\text{latent}}\big\|_{2}^{2}\right)=\frac{{\mathcal{G}}_{t-n-m}^{\text{latent}}-{\mathcal{G}}_{t-n}^{\text{latent}}}{\sigma_{g}^{2}},$
which acts as the gradient of an implicit potential attracting the trajectory toward structure-consistent regions. For each subsequent reverse step $t^{\prime}<k-n$, the guided posterior is defined as
(6) $p_{\theta}({\mathcal{G}}_{t^{\prime}-1}^{\text{latent}}\mid{\mathcal{G}}_{t^{\prime}}^{\text{latent}})=\mathcal{N}\!\Big(\tilde{\mu}_{\theta}({\mathcal{G}}_{t^{\prime}}^{\text{latent}},t^{\prime})+h_{t^{\prime}}\,\tilde{\beta}_{t^{\prime}}\,g_{t-n},\;\tilde{\beta}_{t^{\prime}}\,{\mathbf{I}}\Big),$
where $\tilde{\mu}_{\theta}$ is the standard posterior mean, $\tilde{\beta}_{t^{\prime}}$ is the variance schedule, and $h_{t^{\prime}}$ modulates the strength of structural guidance. Intuitively, the first $n$ latent steps explore freely under unconstrained diffusion, while the subsequent guided updates are softly attracted toward structurally valid manifolds. This allows MCTS to conduct selection and value backup entirely within the latent domain, while periodically anchoring trajectories in the discrete space to enforce validity and stability across long rollouts.
Trajectory Distillation for Auxiliary Modules. Inspired by the success of large language models in distilling high-quality reasoning trajectories, we adopt a similar principle to train the auxiliary components of TreeDiff—namely, the time-conditioned encoder $\mathrm{Enc}(\cdot)$, decoder $\mathrm{Dec}(\cdot)$, and discrete denoiser $p_{\psi}(\cdot)$. Rather than training them jointly with diffusion, we distill knowledge from pre-collected generation trajectories so that these modules capture diffusion-consistent dynamics. Concretely, we generate 500 full denoising trajectories $\{{\mathcal{G}}_{t}\}_{t=0}^{T}$ using a pretrained diffusion model and treat them as teacher demonstrations. The encoder-decoder pair is trained as a time-conditioned variational autoencoder that reconstructs intermediate graphs along these trajectories:
$\mathcal{L}_{\text{VAE}}$ $=\mathbb{E}_{{\mathcal{G}}_{t}}\!\left[\|{\mathcal{G}}_{t}-\mathrm{Dec}(\mathrm{Enc}({\mathcal{G}}_{t},t),t)\|_{2}^{2}\right]$
(7) $\quad+\lambda_{\text{KL}}\,\mathrm{KL}\!\big(q(z_{t}|{\mathcal{G}}_{t},t)\,\|\,\mathcal{N}(0,{\mathbf{I}})\big),$
which encourages temporally smooth and semantically aligned latent representations. Meanwhile, the graph-space denoiser $p_{\psi}$ is trained on consecutive teacher states $({\mathcal{G}}_{t+1}\!\rightarrow\!{\mathcal{G}}_{t})$ to predict the next denoised structure:
(8) $\mathcal{L}_{\text{denoise}}=\mathbb{E}_{t}\!\left[\|\hat{{\mathbf{X}}}_{t}-{\mathbf{X}}_{t}\|_{2}^{2}+\|\hat{{\mathbf{E}}}_{t}-{\mathbf{E}}_{t}\|_{2}^{2}\right],(\hat{{\mathbf{X}}}_{t},\hat{{\mathbf{E}}}_{t})=p_{\psi}({\mathcal{G}}_{t+1},t).$
The architectures of these modules are implementation-agnostic.
4.4. Dual-Space Verifier
In MCTS, the Simulation and Backpropagation phases estimate and propagate the expected reward of each trajectory. Fully rolling out every branch to obtain final graph properties is prohibitively expensive when diffusion involves hundreds of steps and multiple objectives. To address this, TreeDiff introduces a learned dual-space verifier that predicts long-term rewards from partially denoised graphs, replacing costly explicit rollouts. The verifier $V_{\phi}$ is trained on a small set of full trajectories, learning to approximate the terminal value $R({\mathcal{G}}_{0})$ directly from mid-trajectory states. At each reverse step $t$, it takes both the latent representation ${\mathcal{G}}_{t}^{\text{latent}}$ and the decoded discrete structure ${\mathcal{G}}_{t}^{\text{structure}}$ as input and outputs $\hat{R}_{t}\approx R({\mathcal{G}}_{0})$. By jointly reasoning over latent and graph spaces, $V_{\phi}$ remains consistent with dual-space denoising and robust to decoding noise. During search, $\hat{R}_{t}$ replaces full rollouts for scoring and backpropagation, enabling efficient planning under multi-objective settings.
Table 1. Conditional graph generation results on 2D molecules across diversity (Div., %), similarity (Sim., %), mean absolute error (MAE), and accuracy (Acc., %). Boldface and underline indicate the best and second-best, and A.R. is the average ranking.
Synth. & BACE Synth. & BBBP Synth. & HIV
Method Div. $\uparrow$ Sim. $\uparrow$ MAE $\downarrow$ Acc. $\uparrow$ Div. $\uparrow$ Sim. $\uparrow$ MAE $\downarrow$ Acc. $\uparrow$ Div. $\uparrow$ Sim. $\uparrow$ MAE $\downarrow$ Acc. $\uparrow$ A.R.
Tradition
Graph-GA (Jensen, 2019) 85.85 98.05 0.963 46.90 89.50 95.09 1.208 30.15 89.93 96.61 0.984 60.35 7.3
MARS (Xie et al., 2021) 83.38 88.27 1.012 51.84 86.37 76.96 1.225 51.89 87.64 65.17 0.969 64.55 9.5
LSTM-HC (Brown et al., 2019) 81.46 79.82 0.921 58.16 88.83 89.32 0.997 55.90 90.91 91.45 0.948 67.36 7.8
JTVAE-BO (Jin et al., 2018) 66.82 72.81 0.992 46.28 74.58 58.21 1.162 49.58 80.55 41.73 1.236 48.50 11.8
Diffusion
DiGress (Vignac et al., 2023a) 88.62 69.42 2.068 50.61 90.98 68.05 2.366 65.36 91.94 85.62 1.922 53.35 9.1
DiGress v2 (Vignac et al., 2023a) 88.12 70.27 2.337 51.13 91.07 63.36 2.269 65.31 91.93 84.76 1.593 53.31 9.2
GDSS (Jo et al., 2022) 87.56 27.08 1.642 50.36 84.15 26.72 1.379 50.37 78.17 10.32 1.252 48.30 11.9
MOOD (Lee et al., 2023) 89.02 25.87 1.885 50.62 92.73 17.15 2.028 49.03 92.80 13.61 2.314 51.06 9.8
Graph DiT (Liu et al., 2024) 82.38 87.52 0.400 91.35 88.56 93.29 0.355 94.17 89.74 95.75 0.309 97.77 4.8
Inference-Time
Best-of-N (Ma et al., 2025) 87.29 88.50 0.480 91.01 89.29 92.41 0.420 93.80 88.54 95.61 0.360 97.01 5.8
Beam Search (Ma et al., 2025) 86.03 89.21 0.455 91.40 88.53 92.86 0.405 94.11 87.88 95.84 0.345 97.20 5.3
TAGMol (Dorna et al., 2024) 88.64 87.45 0.430 91.60 90.45 91.43 0.372 94.46 89.62 95.17 0.325 97.56 4.5
SVDD (Li et al., 2024) 86.51 89.03 0.510 90.56 88.81 92.69 0.460 93.50 88.04 95.73 0.395 96.83 6.3
TreeDiff (Ours) 89.30 87.82 0.392 91.83 91.62 93.63 0.342 94.81 90.41 96.90 0.302 97.90 1.9
Table 2. Conditional graph generation results on 3D molecular benchmark QM9 (ID & OOD). We report mean absolute error (MAE) on six SE(3)-invariant quantum properties: polarizability $\alpha$ (Bohr3), gap $\Delta\varepsilon$ (meV), HOMO energies $\varepsilon_{H}$ (meV), LUMO energies $\varepsilon_{L}$ (meV), dipole moment $\mu$ (D), and heat capacity $C_{v}$ ($cal/mol\cdot K$).
In-Distribution (MAE $\downarrow$) Out-Of-Distribution (MAE $\downarrow$)
Method $\alpha$(Bohr3) $\Delta\varepsilon$(meV) $\varepsilon_{H}$(meV) $\varepsilon_{L}$(meV) $\mu$(D) $C_{v}$($\frac{cal}{mol}$K) $\alpha$(Bohr3) $\Delta\varepsilon$(meV) $\varepsilon_{H}$(meV) $\varepsilon_{L}$(meV) $\mu$(D) $C_{v}$($\frac{cal}{mol}$K) A.R.
Random 41.00 193.36 103.30 121.83 8.40 13.56 73.03 344.43 183.00 217.01 14.96 24.15 6.3
Diff.
EDM(Wu et al., 2022) 20.15 287.00 158.70 166.20 7.01 13.63 55.70 1561.90 1196.80 228.20 19.13 38.42 7.6
LDM–3DG (You et al., 2024) 15.56 107.14 54.62 63.08 6.33 13.66 32.06 363.13 109.30 178.69 22.18 31.12 4.9
Inference-Time
Best-of-N(Ma et al., 2025) 16.23 109.97 58.11 67.08 6.27 13.54 31.07 359.91 110.13 175.11 14.59 24.07 4.2
Beam Search(Ma et al., 2025) 16.04 108.06 57.14 66.05 6.29 13.58 33.47 345.11 103.23 171.03 14.31 23.79 3.7
TAGMol(Dorna et al., 2024) 15.33 100.19 55.09 63.06 6.13 13.47 31.01 345.08 102.09 171.07 14.77 24.29 2.7
SVDD(Li et al., 2024) 16.78 115.11 60.07 67.39 6.26 13.57 34.47 370.13 115.76 178.11 14.83 23.91 5.2
TreeDiff (Ours) 14.92 95.21 61.55 60.19 6.07 13.41 28.55 341.76 99.53 170.11 14.24 23.51 1.4
The dual-space verifier $V_{\phi}$ is trained as a regression model that predicts the terminal reward $R({\mathcal{G}}_{0})$ from an intermediate state $({\mathcal{G}}_{t}^{\text{latent}},{\mathcal{G}}_{t}^{\text{structure}})$. Training data are obtained from a small set of full denoising rollouts $\{{\mathcal{G}}_{t}\}_{t=0}^{T}$ generated by a pretrained diffusion model, each annotated with the corresponding final reward $R({\mathcal{G}}_{0})=\sum_{k}w_{k}f_{k}({\mathcal{G}}_{0})$ that combines multiple property scores. To improve data diversity and robustness, we further apply a stochastic augmentation procedure: for each latent graph ${\mathcal{G}}_{t}^{\text{latent}}$, we inject Gaussian perturbations $\tilde{{\mathcal{G}}}_{t}^{\text{latent}}={\mathcal{G}}_{t}^{\text{latent}}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma_{a}^{2}{\mathbf{I}})$, and decode $\tilde{{\mathcal{G}}}_{t}^{\text{latent}}$ into $\tilde{{\mathcal{G}}}_{t}^{\text{structure}}$ using $\mathrm{Dec}(\cdot)$. These perturbed pairs $(\tilde{{\mathcal{G}}}_{t}^{\text{latent}},\tilde{{\mathcal{G}}}_{t}^{\text{structure}})$ are included alongside the original rollouts to enrich supervision across local neighborhoods of the latent manifold. The verifier minimizes a mean-squared error objective:
(9) $\mathcal{L}_{\text{verifier}}=\mathbb{E}_{t}\!\left[\|V_{\phi}(\tilde{{\mathcal{G}}}_{t}^{\text{latent}},\tilde{{\mathcal{G}}}_{t}^{\text{structure}},t)-R({\mathcal{G}}_{0})\|_{2}^{2}\right].$
In practice, $V_{\phi}$ adopts a dual-branch architecture: a graph transformer encodes ${\mathcal{G}}_{t}^{\text{structure}}$, a lightweight MLP processes ${\mathcal{G}}_{t}^{\text{latent}}$, and a cross-attention layer fuses the two embeddings before a final projection to the scalar estimate $\hat{R}_{t}$. This design allows the verifier to aggregate global structural patterns and latent diffusion context jointly. During MCTS search, $\hat{R}_{t}$ serves as the node value used for both Simulation and Backpropagation, enabling efficient evaluation of candidate trajectories without full rollouts.
The dual-space verifier closes the search loop in TreeDiff by providing a principled, efficient substitute for full rollouts. Together with macro-step expansion and dual-space denoising, it allows MCTS to operate entirely in latent space while remaining grounded in discrete graph semantics. Moreover, since $R({\mathcal{G}}_{0})$ can represent an arbitrary weighted combination of domain-specific objectives, $V_{\phi}$ naturally supports controllable and multi-objective generation without retraining. This design endows TreeDiff with flexible inference-time reasoning, effectively transforming diffusion sampling into a structured search process guided by learned value estimation.
5. Experiments
5.1. Conditional Graph Generation
Conditional Generation on 2D Molecular Graphs. We evaluate conditional graph generation on three molecular benchmarks, following Liu et al. (2024): BACE (human $\beta$-secretase 1 inhibition), BBBP (blood–brain barrier permeability), and HIV (replication inhibition). Apart from molecular properties, we additionally condition on synthetic accessibility (SAS) and synthetic complexity (SCS) (Ertl and Schuffenhauer, 2009; Coley et al., 2018). Datasets are split into 6:2:26{:}2{:}2 train/validation/test sets (Polykovskiy et al., 2020), and we generate 10,000 molecules per task. Evaluation metrics include diversity (Div., %), fragment-based similarity (Sim., %), mean absolute error (MAE), and classification accuracy (Acc., %). Each experiment is repeated five times and averaged to mitigate randomness. We compare against a broad set of baselines: traditional methods (Graph-GA (Jensen, 2019), MARS (Xie et al., 2021), LSTM-HC (Brown et al., 2019), JTVAE-BO (Jin et al., 2018)); diffusion-based generators (DiGress (Vignac et al., 2023a), DiGress v2, GDSS (Jo et al., 2022), MOOD (Lee et al., 2023), GraphDiT (Liu et al., 2024)); and inference-time guidance strategies, including Best-of-N (Ma et al., 2025), Beam Search (Ma et al., 2025), TAGMol (Dorna et al., 2024), and SVDD (Li et al., 2024). For fairness, all inference-time methods adopt GraphDiT as the backbone, with matched compute budgets. As summarized in Table 1, inference-time guidance consistently improves diffusion baselines, and TreeDiff delivers the most stable and substantial gains across all benchmarks. It achieves the best average rank (A.R. 1.9), outperforming both diffusion and inference-only competitors. In terms of distribution quality, TreeDiff avoids the trade-offs between diversity and similarity, maintaining both at high levels (e.g., Synth.&BACE: Div. 89.30 vs. 82.38 and Sim. 87.82 vs. 87.52 for Graph DiT). For controllability, TreeDiff achieves stronger condition matching, with lower MAE and higher Accuracy across tasks (e.g., Synth.&HIV: MAE 0.302 and Acc. 97.90%).
Table 3. Unconditional graph generation results on 2D QM9 and ZINC250k regarding validity (Valid), Fréchet ChemNet Distance (FCD), NSPDK kernel distance (NSPDK), and scaffold similarity (Scaf.).
QM9 ZINC250k
Method Valid $\uparrow$ FCD $\downarrow$ NSPDK $\downarrow$ Scaf. $\uparrow$ Valid $\uparrow$ FCD $\downarrow$ NSPDK $\downarrow$ Scaf. $\uparrow$ A.R.
Training set 100.00 0.040 0.0001 0.972 100.00 0.062 0.0001 0.840
Trad.
MoFlow (Zang and Wang, 2020) 91.36 4.467 0.0169 0.145 63.11 20.931 0.0455 0.013 10.4
GraphAF (Shi et al., 2020) 74.43 5.625 0.0207 0.305 68.47 16.023 0.0442 0.067 10.0
GraphDF (Luo et al., 2021) 93.88 10.928 0.0636 0.098 90.61 33.546 0.1770 0.000 11.0
Diffusion
EDP-GNN (Niu et al., 2020) 47.52 2.680 0.0046 0.327 87.21 76.737 0.0485 0.000 10.3
GDSS (Jo et al., 2022) 95.72 2.900 0.0038 0.698 97.44 14.656 0.0021 0.047 8.0
DiGress (Vignac et al., 2022) 98.19 0.095 0.0033 0.935 94.99 3.482 0.0021 0.416 5.8
GruM (Jo et al., 2023) 99.69 0.108 0.0002 0.945 98.95 2.257 0.0015 0.530 2.0
Inference-Time
Best-of-N (Ma et al., 2025) 99.22 0.153 0.0003 0.912 98.84 3.803 0.0013 0.485 4.8
Beam Search (Ma et al., 2025) 99.81 0.123 0.0004 0.918 98.12 2.553 0.0015 0.502 3.5
TAGMol (Dorna et al., 2024) 99.63 0.142 0.0004 0.918 98.73 2.647 0.0012 0.488 3.8
SVDD (Li et al., 2024) 99.41 0.205 0.0006 0.905 98.42 2.706 0.0018 0.418 5.9
TreeDiff (Ours) 99.89 0.091 0.0004 0.932 98.95 2.135 0.0011 0.492 1.8
Table 4. Unconditional molecular graph generation results on 3D QM9 and Drugs regarding validity & uniqueness (V&U), atom-level stability (AS), and molecule-level stability (MS).
QM9 Drugs
Method V&U $\uparrow$ AS $\uparrow$ MS $\uparrow$ AS $\uparrow$ A.R.
Trad.
ENF (Garcia Satorras et al., 2021) 39.4 85.0 4.9 16.7
G-SchNet (Gebauer et al., 2019) 80.3 95.7 68.1 15.3
Diffusion
GDM (Hoogeboom et al., 2022) 97.0 63.2 75.0 15.3
GDM-Aug (Hoogeboom et al., 2022) 89.5 97.6 71.6 77.7 12.5
EDM (Wu et al., 2022) 90.7 98.7 82.6 81.3 8.0
EDM-Bridge (Wu et al., 2022) 90.7 98.8 84.6 82.4 6.8
GCDM (Morehead and Cheng, 2024) 93.3 98.7 85.7 89.0 3.8
MiDi (Vignac et al., 2023b) 95.0 97.9 84.0 82.5 7.0
GraphLDM (Xu et al., 2023) 82.7 97.2 70.5 76.2 14.0
GraphLDM-Aug (Xu et al., 2023) 89.5 97.9 78.7 79.6 11.2
GeoLDM (Xu et al., 2023) 92.3 98.2 80.7 79.3 10.2
LDM-3DG (You et al., 2024) 95.3 97.6 87.7 80.51 6.2
Inference-Time
Best-of-N (Ma et al., 2025) 93.9 98.6 87.5 85.4 4.0
Beam Search (Ma et al., 2025) 93.2 98.5 87.1 84.6 5.5
TAGMol (Dorna et al., 2024) 92.9 98.7 85.5 85.1 5.2
SVDD  (Li et al., 2024) 92.8 98.4 86.3 84.2 6.8
TreeDiff (Ours) 94.8 98.8 88.6 86.5 1.8
[…]
5.2. Unconditional Graph Generation
Unconditional Generation on 2D Molecular Graphs. To assess model performance without explicit conditional supervision in training, we evaluate unconditional molecular graph generation on QM9 (Ramakrishnan et al., 2014) and ZINC250k (Irwin et al., 2012). Following standard practice, we report Validity (chemical sanity without rule-based repair), distributional fidelity via Fréchet ChemNet Distance (FCD) (Preuer et al., 2018), global structure mismatch via the NSPDK kernel distance (Costa and De Grave, 2010), and Scaffold similarity. These metrics also serve as guidance signals for inference-time baselines. We compare against traditional methods (MoFlow (Zang and Wang, 2020), GraphAF (Shi et al., 2020), GraphDF (Luo et al., 2021)), diffusion-based generators (EDP-GNN (Niu et al., 2020), GDSS (Jo et al., 2022), DiGress (Vignac et al., 2022), GruM (Jo et al., 2023)), and inference-time guidance strategies (Best-of-$N$ (Ma et al., 2025), Beam Search (Ma et al., 2025), TAGMol (Dorna et al., 2024), SVDD (Li et al., 2024)). We also report the training set as a reference upper bound for distributional metrics. As shown in Table 3, TreeDiff achieves the lowest FCD on both datasets and near-perfect validity, while remaining competitive on NSPDK and scaffold similarity. Compared with both diffusion backbones and inference-time guidance methods, TreeDiff provides a more reliable balance between distributional fidelity and chemical plausibility. These results highlight the effectiveness of planning-guided sampling in advancing unconditional molecular graph generation.
Unconditional Generation on 3D Molecular Graphs. We further evaluate TreeDiff on unconditional 3D molecular generation using QM9 (Ramakrishnan et al., 2014) and Drugs (Axelrod and Gomez-Bombarelli, 2022) as benchmarks. Following established protocols (Hoogeboom et al., 2022; Wu et al., 2022; Vignac et al., 2023b; Morehead and Cheng, 2024), we report validity & uniqueness (V&U), atom-level stability (AS), and molecule-level stability (MS) on QM9, and AS on Drugs. Baselines span a broad range of 3D generators, including traditional methods (ENF (Garcia Satorras et al., 2021), G-SchNet (Gebauer et al., 2019)); diffusion-based models (GDM (Hoogeboom et al., 2022), GDM-Aug, EDM (Wu et al., 2022), EDM-Bridge, GCDM (Morehead and Cheng, 2024), MiDi (Vignac et al., 2023b), GraphLDM (Xu et al., 2023), GraphLDM-Aug, GeoLDM, LDM-3DG (You et al., 2024)); and inference-time strategies (Best-of-$N$ (Ma et al., 2025), Beam Search, TAGMol (Dorna et al., 2024), SVDD (Li et al., 2024)). As shown in Table 4, TreeDiff achieves the best overall performance on QM9 and the second-best AS on Drugs, yielding the lowest overall average rank (A.R. 1.8) across benchmarks. On QM9, it attains the highest molecule-level stability (MS 88.6) and state-of-the-art atom stability tied with EDM-Bridge (AS 98.8) while maintaining strong validity and uniqueness (94.8). Compared with diffusion backbones (e.g., LDM-3DG, GCDM), TreeDiff substantially reduces whole-molecule failures through global planning. Against inference-time methods, it preserves their high atom-level reliability but produces more globally consistent structures. On Drugs, TreeDiff maintains strong atom stability (86.5), confirming its robustness on larger, drug-like molecules. Overall, these results highlight TreeDiff’s ability to couple local atomic precision with global structural coherence in 3D molecular generation.
5.3. Ablation Study
Table 5. Ablation results of the dual-space denoising on 2D/3D unconditional generation tasks, where $\sigma$ is used to control the impact of the dual-space denoising, where a large value indicate less impact.
2D QM9 3D QM9
Variant Valid $\uparrow$ FCD $\downarrow$ NSPDK $\downarrow$ V&U $\uparrow$ AS $\uparrow$ MS $\uparrow$
TreeDiff ($\mathbf{\sigma=1.0}$) 99.90 0.090 0.00040 94.8 98.8 88.6
$\mathbf{\sigma=0.1}$ 97.56 0.256 0.00100 92.5 98.6 86.9
$\mathbf{\sigma=0.5}$ 99.48 0.112 0.00055 93.4 98.7 87.4
$\mathbf{\sigma=10}$ 99.15 0.125 0.00064 92.4 98.5 86.8
w/o graph space 98.13 0.156 0.00087 92.1 97.2 86.8
Table 6. Ablation results on the verifier on 2D/3D conditional generation, where Latent + GPM is the default.
BBBP QM9 (ID) QM9 (OOD)
Verifier Type Acc $\uparrow$ MAE $\downarrow$ $\alpha$ $\downarrow$ $\varepsilon_{H}$ $\downarrow$ $\alpha$ $\downarrow$ $\varepsilon_{H}$ $\downarrow$
Latent-only 92.4 0.405 17.2 69.5 36.0 116.2
Graph-only (GIN (Xu et al., 2019)) 92.8 0.395 16.6 68.3 34.8 113.4
Graph-only (GT (Rampášek et al., 2022)) 93.3 0.380 15.8 64.5 32.5 106.5
Graph-only (GPM (Wang et al., 2025a)) 93.8 0.381 15.6 65.3 33.6 104.1
Latent + GIN (Xu et al., 2019) 93.8 0.374 16.1 66.8 33.5 110.7
Latent + GT (Rampášek et al., 2022) 94.5 0.350 15.3 63.0 31.0 103.5
Latent + GPM (Wang et al., 2025a) 94.8 0.342 14.9 61.6 28.6 99.5
Ablation on dual-space denoising. We investigate the impact of dual-space denoising strength by varying the guidance scale $\sigma\!\in\!\{0.1,0.5,1.0,10\}$ on 2D and 3D QM9 datasets under unconditional settings. A w/o graph space variant is also included as a latent diffusion-only baseline. As shown in Table 5, excessively strong guidance ($\sigma{=}0.1$) over-constrains the denoising trajectory, leading to degraded validity and fidelity, whereas overly weak guidance ($\sigma{=}10$) approaches standard diffusion and offers limited improvement. The default configuration ($\sigma{=}1.0$) achieves the best overall performance across both 2D and 3D metrics, demonstrating that moderate dual-space coupling substantially enhances generation quality compared to the diffusion-only baseline.
Ablation on the verifier. We examine the impact of verifier design under identical computational budgets on conditional generation tasks across BBBP, QM9 (ID), and QM9 (OOD). We consider three categories of verifiers: Latent-only, Graph-only (GIN, Graph Transformer (GT), and GPM), and Latent+Graph (Latent+GIN, Latent+GT, Latent+GPM). As summarized in Table 6, the Latent-only verifier provides fast feedback but lacks discrete structural cues, resulting in weaker calibration. Graph-only variants better capture structural constraints: GIN leverages local substructure sensitivity, while GT reduces prediction errors through long-range aggregation. Combining latent and graph modalities consistently improves performance—both Latent+GIN and Latent+GT surpass their single-modality counterparts. Our default Latent+GPM achieves the strongest results overall, indicating that global attention enhanced by graph-structural priors yields more accurate and stable node-wise calibration across noise scales.
5.4. Efficiency Analysis
Table 7. Efficiency analysis on the verifier for 2D/3D conditional generation.
BACE QM9 (ID)
Variant Acc $\uparrow$ MAE $\downarrow$ Comp. $\varepsilon_{L}$ $\downarrow$ $C_{v}$ $\downarrow$ Comp.
Post-hoc + TreeDiff Verify 87.1 0.409 9.6x 59.5 13.0 15.6x
Post-hoc + Oracle Verify 86.7 0.420 8.9x 60.0 13.5 14.2x
Step-wise + TreeDiff Verify 87.8 0.392 1.0x 60.2 13.4 1.0x
Efficiency analysis on the verifier. We compare Post-hoc verification, which evaluates only the final denoising step (t=0t{=}0), against two baselines—Post-hoc + Oracle Verify and Post-hoc + TreeDiff Verify—as well as our proposed Step-wise + TreeDiff Verify. As shown in Table 7, step-wise verification achieves the best overall balance between accuracy and efficiency, reducing computation cost by over $9\times$ on BACE and $15\times$ on QM9 (ID) while maintaining comparable or even slightly better predictive accuracy. In contrast, post-hoc variants require substantially higher compute since verification occurs only after full trajectory generation.
Figure 4. Efficiency analysis on the MCTS tree depth vs. inference-time computation.
Efficiency analysis on tree depth. We vary the macro-step budget (tree depth) $D\!\in\!\{1,5,10,50,1000\}$ and report task performance alongside inference compute (NFE) for two conditional generation settings: 2D BBBP and 3D QM9 (ID). As reported in Figure 4, model performance improves rapidly for small depths between D=1D{=}1 and D=10D{=}10 then plateaus beyond $D{\geq}50$. Meanwhile, NFE grows near-geometrically with $D$ due to repeated expansions and evaluations, revealing a clear compute-quality trade-off. A practical Pareto regime emerges at $D{\in}[10,50]$, which captures most quality gains while keeping computation moderate.
6. Conclusion
We propose TreeDiff, an inference-time framework that couples MCTS with dual-space diffusion to enable controllable and stable graph generation without retraining. The architecture comprises three key components: (i) macro-step expansion that operates over groups of denoising steps to reduce search depth, (ii) dual-space denoising that jointly leverages latent efficiency and graph-level structural fidelity, and (iii) a dual-space verifier that provides rollout-free node-wise evaluation for early pruning. Extensive experiments on 2D/3D molecular generation, across unconditional and conditional regimes and ID/OOD settings, demonstrate the state-of-the-art performance and favorable inference-time scaling of TreeDiff.
Appendix A Pseudo Code
TreeDiff’s inference procedure is summarized in Algorithm 1.
Algorithm 1 TreeDiff Inference
1:Diffusion predictor $\epsilon_{\theta}$; total steps $T$; start step $t_{s}$; encoder/decoder $(\mathrm{Enc},\mathrm{Dec})$; discrete refiner $p_{\psi}$; verifier $V_{\phi}$; max children $K$; rollouts per round $N_{r}$; depth budget $D_{\max}$; UCB constant $c_{\mathrm{ucb}}$.
2:Final graph $\mathcal{G}_{0}$.
3:Initialize: latent root $z_{t_{s}}$; set node $\mathcal{R}\!\leftarrow\!z_{t_{s}}$ with stats Q=0Q{=}0, N=0N{=}0; set remaining depth $D{\leftarrow}D_{\max}$.
4:while $t_{s}>0$ do
5:  for $i=1$ to $N_{r}$ do
6:   Selection: From node $u$ (latent state $z_{t}$), descend by UCB until a leaf/terminal.
7:   Expansion: At node $u$ with time $t$, propose up to $K$ children via:
  • Jump (macro-step): sample a step length $k$ (e.g., around $\lfloor t/D\rfloor$) and advance $z_{t}\!\to\!z_{t-k}$ by $k$ reverse latent updates.
  • Dual-Space Denoising (Sec. Dual-Space Denoising): (a) decode $\mathcal{G}_{t-k}^{\text{str}}\!=\!\mathrm{Dec}(z_{t-k},\,t{-}k)$; (b) apply lightweight discrete refinement with $p_{\psi}$; (c) re-encode to obtain a structure-aware latent $\tilde{z}_{t-k}$ to anchor the child state.
8:   The resulting child stores both latent $\tilde{z}_{t-k}$ and its decoded graph $\tilde{\mathcal{G}}_{t-k}^{\text{str}}$.
9:   Simulation (Sec. Dual-Space Verifier): For each child, score with $V_{\phi}\big(\tilde{z}_{t-k},\,\tilde{\mathcal{G}}_{t-k}^{\text{str}},\,t{-}k\big)$; initialize $Q(\cdot)$ and $N(\cdot)$ accordingly.
10:   Backpropagation: Propagate the child score along the selected path (update visits and values).
11:  end for
12:  Commit: Select the best child of the root by average value; set it as new root $\mathcal{R}$. Update $t_{s}\!\leftarrow\!t_{s}-k(\mathcal{R})$, $D\!\leftarrow\!\max(D{-}1,1)$. (Optionally prune siblings to keep top-$M$.)
13:end while
14:Return: $\mathcal{G}_{0}\leftarrow\mathrm{Dec}(z_{0},\,0)$.
Procedure. Starting from a latent root $z_{t_{s}}$, TreeDiff performs a sequence of short MCTS rounds to construct a single high-quality denoising path. In each round, Selection first descends from the current root to a leaf via UCB, focusing computation on promising branches. Given the selected leaf at time $t$, Expansion then proposes up to $K$ children in two phases: a macro-step jump that advances $k$ reverse updates in latent space, followed by dual-space denoising—decode to graph space, apply lightweight discrete refinement, and re-encode—to anchor a structure-aware latent state. Next, Simulation evaluates each child using the dual-space verifier on both the latent and its decoded graph, avoiding costly full rollouts while preserving structural feedback. Finally, Backpropagation updates visit counts and value estimates along the traversed path. After accumulating $N_{r}$ rollouts, TreeDiff commits the best child as the new root, reduces $t$ by $k$, and repeats until t=0t{=}0; the resulting $z_{0}$ is then decoded to $\mathcal{G}_{0}$.
Appendix B Time Complexity
Let $C_{\text{lat}}$ be the cost of one latent reverse step, $C_{\text{ref}}$ the cost of one dual-space refinement (decode ++ discrete refine ++ re-encode), and $C_{\text{ver}}$ the cost of one verifier call. At each visited node, UCB-based Selection scans its children; if the branching factor is bounded by $K$, the path cost is $O(K\,D_{\max})$ (or $O(D_{\max})$ when child stats are cached and $K$ is treated as a constant).
For a leaf that is expanded into up to $K$ children, the per-child cost of Expansion++Simulation is
$O\!\big(k\,C_{\text{lat}}+C_{\text{ref}}+C_{\text{ver}}\big),$
where $k$ is the sampled macro-step length for that child. Therefore, one MCTS round with $N_{r}$ rollouts incurs
$O\!\Big(N_{r}\big[K\,D_{\max}\;+\;K\big(k\,C_{\text{lat}}+C_{\text{ref}}+C_{\text{ver}}\big)\big]\Big).$
Across the whole inference, the number of commits is roughly $T/\bar{k}$ (with $\bar{k}$ the average macro-step), and never exceeds $D_{\max}$. Using $\bar{k}$ to summarize $k$’s variability, the total work is
$O\!\Big(\frac{T}{\bar{k}}\;N_{r}\big[K\,D_{\max}\;+\;K\big(\bar{k}\,C_{\text{lat}}+C_{\text{ref}}+C_{\text{ver}}\big)\big]\Big).$
If $K$ is constant and selection bookkeeping is $O(1)$ per level, this simplifies to
$O\!\Big(\frac{T}{\bar{k}}\;N_{r}\big[D_{\max}\;+\;\bar{k}\,C_{\text{lat}}+C_{\text{ref}}+C_{\text{ver}}\big]\Big).$
Discussion.(i) Trade-off via $\bar{k}$. Larger $\bar{k}$ reduces the number of commit steps ($T/\bar{k}$) but increases the per-child expansion cost ($\bar{k}C_{\text{lat}}$); small $\bar{k}$ does the opposite. (ii) Verifier amortization. When $V_{\phi}$ is a lightweight MLP, $C_{\text{ver}}\ll\bar{k}C_{\text{lat}}$, so runtime is dominated by latent steps and occasional refinement. If $V_{\phi}$ batches $B$ children on accelerator, its amortized cost becomes $C_{\text{ver}}/B$. (iii) Parallelism. The $K$ children of a leaf and the $N_{r}$ rollouts within a round are embarrassingly parallel; with $P$ workers the effective factor becomes $K/P$ or $N_{r}/P$ in the bracketed term. (iv) Degenerate cases. With K=1K{=}1 and $N_{r}{=}1$, TreeDiff reduces to a single guided trajectory with cost $O\!\big(T\,C_{\text{lat}}+\tfrac{T}{\bar{k}}(C_{\text{ref}}{+}C_{\text{ver}})\big)$. If dual-space refinement is skipped ($C_{\text{ref}}{=}0$), complexity coincides with pure latent MCTS. (v) Memory. Active-node storage is $O(N_{r}K)$ per round without pruning, or $O(MD_{\max})$ when keeping only top-$M$ siblings per level.
Appendix C Implementation Details
All diffusion backbones are pretrained under standard noise schedules before integration into TreeDiff. For macro-step expansion, we set the tree depth to D=10D{=}10 and expansion width to B=4B{=}4, with an adaptive expansion variance $\sigma_{k}{=}0.1$, and use PUCT as the selection policy. For dual-space denoising, the guidance scale is $\sigma_{g}{=}1.0$ and $h_{t^{\prime}}{=}1.0$; the latent denoising horizon is set to half of the total diffusion steps within each macro step (n=k/2n{=}k/2), while the structural refinement length is set to 10% of that value (m=k/10m{=}k/10). The encoder, decoder, and verifier are trained through trajectory distillation and remain frozen during search. For 2D molecular generation, we employ a Graph Transformer (Rampášek et al., 2022) as the encoder–decoder backbone, and for 3D molecular generation, we use a SE(3)-Transformer (Fuchs et al., 2020). The verifier adopts the recently proposed GPM (Wang et al., 2025a), which excels at capturing substructural dependencies in graphs.